OpenStudy (anonymous):

cot Θ=2.127 I know the answer is 25.2 degrees, I just need to know how to do it

OpenStudy (mertsj):

Change it to its reciprocal. That will be the tangent. Then push the inverse tan button on your calculator.

OpenStudy (mertsj):

\[\cot \theta=2.127, \tan \theta=\frac{1}{2.127}\]

OpenStudy (anonymous):

Thanks!! it was really simple haha.

OpenStudy (mertsj):

\[\theta=\tan^{-1} (.470146)= 25.2\]

OpenStudy (mertsj):

yep. Most things are when you understand them.